OpenStudy (anonymous):

A 10 ft. tree casts a 14 ft. shadow. If a building casts a 280 ft. shadow, how tall is the building? will fan

OpenStudy (anonymous):

200

OpenStudy (anonymous):

112

OpenStudy (anonymous):

which one is it 112 or 200?

OpenStudy (anonymous):

200

OpenStudy (anonymous):

112

OpenStudy (anonymous):

@codyisbeast

OpenStudy (anonymous):

280*10=2800 2800/14=200

OpenStudy (anonymous):

what 40% of 10 4 so the shadow is 40% bigger

OpenStudy (janu16):

Its 200 for sure

OpenStudy (anonymous):

\[\frac{ 10 }{ 14 } =\frac{ x }{ 240 }\]

OpenStudy (janu16):

@ironsnow 200 for sure

OpenStudy (anonymous):

now minse 40% of 280 is 112

OpenStudy (anonymous):

thank you janu16

OpenStudy (anonymous):

oops nvm

OpenStudy (anonymous):

Ayyee